# Non-perturbative renormalization group for Higgs-like models in 4D

### v2

@@ II Definition of the models
III Projecting onto positive norm states and unitarity
Our models could have applications to statistical mechanics in 4 spatial dimensions where the positivity constraints of unitarity are not so stringent. However, for potential applications to real Higgs physics in $3+1$ spacetime dimensions, a consistent quantum mechanics would be desirable, and this is the subject of this section. The subject of Quantum Mechanics for pseudo-hermitian hamiltonians is by now well developed, with extensive reviews Mosta1 ; Mosta2 ; Mosta3 . In this section we summarize the main properties as they specialize to our model.
(i)   Hamiltonian and Hilbert space
Assume we are given a Hilbert space ${\cal H}$ with a conventional positive definite inner product. Namely for $|\psi\rangle\in{\cal H}$, there exists an inner product such that $\langle\psi|\psi\rangle$ is positive definite. At time $t=0$ the states need to be properly normalized such that $\langle\psi|\psi\rangle=1$. There exists a basis in ${\cal H}$, $|\psi_{n}\rangle\in{\cal H}$ where for simplicity of notation we assume $n$ is discrete, $n=1,2,\ldots\infty$. Then
$|\psi\rangle=\sum_{n}c_{n}\,|\psi_{n}\rangle,~{}~{}~{}~{}\langle\psi_{m}|\psi_{n}\rangle=\delta_{nm},$ (24)
for some complex numbers $c_{n}$. Then
$\langle\psi|\psi\rangle=1,~{}~{}~{}~{}~{}~{}\Longrightarrow~{}~{}~{}\sum_{n}c_{n}^{*}c_{n}=1,$ (25)
where $c_{n}^{*}$ denotes complex conjugation. $|c_{n}|^{2}$ thus represents the probability that $|\psi\rangle$ is measured to be in the state $|\psi_{n}\rangle$. Given a hamiltonian $H$, or any other operator $A$ on ${\cal H}$, then $A^{\dagger}$ is defined in the standard way with respect to the above inner product.
(ii)   Modified metric for pseudo-hermitian Hamiltonians
Suppose the hamiltonian operator $H$ is pseudo-hermitian with the structure in (1), which we repeat here:
This leads us to define a new inner product which includes an insertion of the operator ${\cal K}$:
$\langle\!\langle\psi^{\prime}|\psi\rangle\!\rangle\equiv\langle\psi^{\prime}|{\cal K}|\psi\rangle.$ (27)
In simple words, kets have an accompanying ${\cal K}$ but bras do not, or visa-versa since ${\cal K}^{\dagger}={\cal K}$. In general, this new inner product has negative norm states. Below we will show that such states can be consistently projected out.
(iii)   Pseudo-hermitian operators correspond to observables with real eigenvalues
For any operator $A$ on ${\cal H}$, define its pseudo-hermitian conjugate as follows:
Then the hamiltonian is pseudo-hermitian:
$H^{\dagger_{\scriptscriptstyle\kappa}}=H.$ (29)
$A^{\dagger_{\scriptscriptstyle\kappa}}$ is the proper conjugation based on the ${\cal K}$-metric in (27), namely
where * denotes ordinary complex conjugation. We define a pseudo-hermitian operator $A$ as one that satisfies $A^{\dagger_{\scriptscriptstyle\kappa}}=A$. From (30), one concludes that pseudo-hermitian operators, in particular the hamiltonian H, has real eigenvalues. Any operator satisfying $A^{\dagger_{\scriptscriptstyle\kappa}}=A$ in principle can correspond to an observable. One can easily establish that this pseudo-hermitian adjoint satisfies the usual rules, e.g.
$(AB)^{\dagger_{\scriptscriptstyle\kappa}}$ $=$ $B^{\dagger_{\scriptscriptstyle\kappa}}A^{\dagger_{\scriptscriptstyle\kappa}}\,,$
$(aA+bB)^{\dagger_{\scriptscriptstyle\kappa}}$ $=$ $a^{*}A^{\dagger_{\scriptscriptstyle\kappa}}+b^{*}B^{\dagger_{\scriptscriptstyle\kappa}}\,,$ (31)
where $A,B$ are operators and $a,b$ are complex numbers.
(iv)   Conservation of probabilities
The hamiltonian determines the time evolution of a state $|\psi\rangle=|\psi(t=0)\rangle$, namely $|\psi(t)\rangle=U(t)\,|\psi\rangle$, where $U(t)=e^{-iHt}$. By conservation of probability we mean that the norm of states is preserved under time evolution. More generally:
$\langle\!\langle\psi^{\prime}(t)|\psi(t)\rangle\!\rangle=\langle\psi^{\prime}|e^{iH^{\dagger}t}\,{\cal K}\,e^{-iHt}\,|\psi\rangle=\langle\psi^{\prime}|{\cal K}e^{iH^{\dagger}t}{\cal K}^{2}\,e^{-iHt}\,|\psi\rangle=\langle\!\langle\psi^{\prime}(0)|\psi(0)\rangle\!\rangle$ (32)
In general, the modified ${\cal K}$-metric $\langle\!\langle\psi^{\prime}|\psi\rangle\!\rangle$ introduces negative norm states which imply negative probabilities. The above statement of unitarity just states that the sum of these probabilities is maintained under time evolution, but one must still deal with negative norm states which implies negative probabilities.
(v)   Projection onto positive norm states
For the main example of this paper, the ${\cal K}$-metric on the Hilbert space is indefinite, namely there are negative norm states which naively implies negative probabilities, which are clearly unphysical. However the properties of ${\cal K}$ naturally leads to projection operators which can fix this problem. Since ${\cal K}^{2}=1$, states can be classified according to ${\cal K}=\pm 1$. Namely,
${\cal K}|\psi_{n}\rangle=(-1)^{k_{n}}\,|\psi_{n}\rangle,~{}~{}~{}~{}k_{n}\in\{0,1\}.$ (33)
In fixing the norm of a state to be unity, the sum involves negative terms, which would be interpreted as negative probabilities:
$\langle\!\langle\psi|\psi\rangle\!\rangle=\sum_{n}(-1)^{k_{n}}\,|c_{n}|^{2}=1.$ (34)
Define the projectors
$P_{\pm}=\dfrac{1\pm{\cal K}}{2}.$ (35)
They satisfy the usual algebra of projectors:
$P_{+}+P_{-}=1,~{}~{}~{}~{}~{}P_{\pm}^{2}=P_{\pm},~{}~{}~{}~{}~{}P_{+}P_{-}=0.$ (36)
In addition, one has
${\cal K}P_{\pm}=P_{\pm}K=\pm P_{\pm}.$ (37)
Thus the original Hilbert space decomposes as
${\cal H}={\cal H}_{+}\oplus{\cal H}_{-},~{}~{}~{}~{}~{}|\psi_{\pm}\rangle\in{\cal H}_{\pm},~{}~{}~{}~{}~{}{\cal K}|\psi_{\pm}\rangle=\pm|\psi_{\pm}\rangle.$ (38)
One has
$\langle\!\langle\psi_{\pm}|\psi_{\pm}\rangle\!\rangle=\langle\psi|P_{\pm}{\cal K}P_{\pm}|\psi\rangle=\pm\langle\psi_{\pm}|\psi_{\pm}\rangle,~{}~{}~{}~{}~{}\langle\!\langle\psi_{-}|\psi_{+}\rangle\!\rangle=\langle\psi|P_{+}{\cal K}P_{-}|\psi\rangle=0$ (39)
If we project onto the sub Hilbert space ${\cal H}_{+}$, then all states $|\psi_{n}\rangle\in{\cal H}_{+}$ have positive norm, i.e. $k_{n}=0$, which implies positive probabilities $|c_{n}|^{2}\geq 0$. This positivity is preserved under time evolution:
$\langle\!\langle\psi_{\pm}(t)|\psi_{\pm}(t)\rangle\!\rangle=\langle\!\langle\psi_{\pm}(0)|\psi_{\pm}(0)\rangle\!\rangle,~{}~{}~{}~{}~{}\langle\!\langle\psi_{-}(t)|\psi_{+}(t)\rangle\!\rangle=\langle\!\langle\psi_{-}(0)|\psi_{+}(0)\rangle\!\rangle=0.$ (40)
In general there could be operators $A$ on ${\cal H}$ that mix ${\cal H}_{+}$ and ${\cal H}_{-}$. If one projects ${\cal H}$ onto ${\cal H}_{+}$, then one should restrict to operators which when acting on $|\psi_{+}\rangle$ yields a state that is also in ${\cal H}_{+}$. In other words
$\langle\!\langle\psi_{-}|A|\psi_{+}\rangle\!\rangle=0.$ (41)
It is easy to see that this requires $A$ to be ${\cal K}$-invariant:
$A={\cal K}A{\cal K}.$ (42)
The proof is simple:
$\langle\!\langle\psi_{-}|A|\psi_{+}\rangle\!\rangle=\langle\psi_{-}|{\cal K}A|\psi_{+}\rangle=-\langle\psi_{-}|A|\psi_{+}\rangle=\langle\psi_{-}|A{\cal K}|\psi_{+}\rangle=+\langle\psi_{-}|A|\psi_{+}\rangle,$ (43)
which implies (41). The condition (42) implies $A^{\dagger}={\cal K}A^{\dagger}{\cal K}$. Thus $A^{\dagger_{\scriptscriptstyle\kappa}}=A$ requires $A^{\dagger}=A$.
In summary, if we project onto the Hilbert space ${\cal H}_{+}$, then all probabilities are positive, and remain positive under time evolution. Let us now turn to the specifics of our model. The Hilbert space ${\cal H}$ has a basis consisting of n-particle states with specific momenta:
$|\psi\rangle=|({\bf{k}}_{1},s_{1}),({\bf{k}}_{2},s_{2}),\ldots({\bf{k}}_{n},s_{n})\rangle\equiv a^{\dagger}_{s_{1}}({\bf{k}}_{1})a^{\dagger}_{s_{2}}({\bf{k}}_{2})\cdots a^{\dagger}_{s_{n}}({\bf{k}}_{n})|0\rangle~{}~{}~{}~{}~{}s_{i}\in\{\pm\}.$ (44)
Mapping the labels $s_{i}=\pm$ to the integers $s_{i}=\pm 1$, define $\widehat{s}=(1-s)/2\in\{0,1\}$. Then
${\cal K}\,|\psi\rangle=(-1)^{\sum_{i}\widehat{s}_{i}}\,\,|\psi\rangle.$ (45)
This simple action on momentum space acts non-locally on the fields $\Phi(x)$ in Section II. In short, ${\cal K}=1$ for states with an even number of $a^{\dagger}_{-}({\bf{k}})$, whereas ${\cal K}=-1$ for an odd number. Projection onto states of positive norm with ${\cal K}=1$ leads to the Hilbert space ${\cal H}_{+}$ where states created by $a^{\dagger}_{-}$ must come in pairs, reminiscent of Cooper pairs.
Above we interpreted the operators $a_{\pm}^{\dagger}$ as creating ${\cal P}{\cal T}$ even/odd states. Thus projecting onto ${\cal H}_{+}$ breaks ${\cal P}{\cal T}$ symmetry.
IV Higher order RG beta-functions from the OPE
@@
In this section we compute the beta functions to 3-loops for the fully anisotropic case $g_{1}\neq g_{2}\neq g_{3}$ based only on the OPE (18). We are more interested in the case where SU(2) is broken to U(1), where $g_{1}=g_{2}\neq g_{3}$ considered in the next section, however it’s easier to work out the beta functions for the fully anisotropic case first and then set $g_{1}=g_{2}$. Renormalization group beta-functions are generally prescription, or scheme, dependent. However for marginal perturbations, the one and two loop contributions are universal. [^5] To see this, consider a single coupling $g$ with $\beta_{g}=b_{2}g^{2}+b_{3}g^{3}+\ldots$ where $b_{2,3}$ are coefficients. The prescription dependence corresponds to a redefinition of the coupling $g^{\prime}=g^{\prime}(g)$. Let $g^{\prime}=g+cg^{2}+\ldots$. One easily sees that $\beta^{\prime}(g^{\prime})=b_{2}g^{\prime 2}+b_{3}g^{\prime 3}+\ldots.$ The implicit prescription of this article is based on the OPE (18), and doesn’t rely on any epsilon-expansion, nor Feynman diagrams for scalar fields. [^6] We do not address whether this OPE prescription can be shown to be equivalent to Feynman diagram perturbation theory and dimensional regularization. Referring to the title of Wilson-Fisher’s paper WilsonFisher , we work in precisely $4$ dimensions, not $3.99$!
@@
where $\langle X\rangle_{0}$ is the unperturbed CFT result. The $n$-th term gives a contribution to the $(n-1)$-loop $\beta$ function.
IV.1 1-loop
@@
IV.2 2-loops
As we did above at 1-loop, at higher orders we use the OPE prescription to isolate $\log a$ divergences that contribute to the beta-functions. There are also $(\log a)^{n}$ divergences for $n>1$, which we interpret as arising from lower orders in perturbation theory which are already accounted for in the lower order beta-function. It turns out that the OPE (47) is insufficient to extract the appropriate $\log a$ contributions and one must split ${\cal O}^{a}$ into the product of $J^{a}$ and $\widetilde{J}^{a}$. It is useful to introduce diagrams that indicate the OPE structure that leads to a $\log a$ divergence of interest, with the building blocks shown in in Figure 1.777Unlike Feynman diagrams, these diagrams are not a prescription for calculating the full correlation function, but are useful for isolating $\log a$ divergences. The 1-loop contribution is shown in Figure 2.
@@
The above, based on Figure 3, is the only contribution at order $g^{3}$. For instance the diagram in Figure 4 gives a $\log^{2}a$ divergence which is already incorporated in the 1-loop beta function. The diagram in Figure 5 gives zero since it is proportional to $\delta^{ab}f^{abc}$. Figure 6 diagram could potentially lead to new terms $J^{a}(x)J^{a}(x)$ in the action rendering our original model non-renormalizable, however as is evident from the diagram, the two $J$’s are at different spacetime points, and furthermore leads to $\log^{2}a$ divergences.
@@
Figure 6: 2-loop diagram that potentially gives rise to additional $JJ$ terms in the action.
IV.3 3 loops
@@
$\beta_{g_{1}}=g_{2}g_{3}-\dfrac{\kappa}{4}g_{1}\left(g_{2}^{2}+g_{3}^{2}\right)+\dfrac{\kappa^{2}}{16}\left(g_{1}^{2}g_{2}g_{3}+g_{2}^{3}g_{3}+g_{3}^{3}g_{2}\right)+\ldots$ (62)
plus cyclic permutations of $1,2,3$.
@@
V SU(2) broken to U(1) and higher orders
V.1 3 loops
@@ V.2 Conjecture for all orders
In order to sum these higher order contributions, we compare with the current-current perturbations in 2D reviewed in the Appendix. Based on this we propose the following beta functions:
@@
One can easily check that to order $g^{4}$, the above agree exactly with the 3-loop results (64). This is unexpected since the integrals involved are quite different. These beta functions preserve the RG invariant $Q$, which to all orders is now:
@@
The invariance condition (66) is exactly satisfied to all orders. The above beta functions do not constitute a gradient flow where $\beta_{g_{i}}=-\partial_{g_{i}}{\cal H}(g)$ for some height function ${\cal H}(g)$, since $\partial_{g_{1}}\beta_{g_{3}}\neq\partial_{g_{3}}\beta_{g_{1}}$. This is consistent with the existence of cyclic RG flows we find below since gradient flow is often associated with flows that are consistent with various forms of c-theorems.
@@
VI Renormalization group trajectories
[…]
@@ VI Renormalization group trajectories
(i) $g_{3}>0$. Here the perturbation is relevant with $40<\Gamma<$ 4.
(ii) $-4<g_{3}<0$. These are marginally irrelevant perturbations with $\Gamma>4$.
@@
$\beta_{\widetilde{g}}(\widetilde{g})\equiv\dfrac{\partial\widetilde{g}}{dg}\,\beta_{g}=-\beta_{g}(g\to\widetilde{g}).$ (75)
@@
It was also shown in BLflow that it is consistent, and required, to endow the coupling constant space $(g_{1},g_{3})$ with the topology of a cylinder where one identifies $g_{3}=\pm\infty$, since $\beta(g_{1},g_{3})=\beta(g_{1},-g_{3})$ in the limit $|g_{3}|\to\infty$. This allows us to extend the flows to the entire $(g_{1},g_{3})$ plane, including through the poles at $g_{1}=\pm 4$, which are self-dual points under (74). We refer the reader to BLflow for more detailed explanations. This leads a variety of RG flows which we now itemize.
VI.1 Massive Flows with fixed points
For case (i) above, namely $g_{3}>0$, the perturbation is relevant $40<\Gamma<$ 4 and flows to strong coupling in the IR. We interpret this as a massive phase where the IR fixed point is empty. In Figure 9 we plot the constant $Q$ RG trajectories based only on the 3 loop result (67), and one sees that the 1-loop flowchart is not significantly modified. In comparison with the current-current perturbations reviewed in the Appendix, this is a sine-Gordon like phase.

### v3

@@ II Definition of the models
III Higher order RG beta-functions from the OPE
@@
In this section we compute the beta functions to 3-loops for the fully anisotropic case $g_{1}\neq g_{2}\neq g_{3}$ based only on the OPE (23). We are more interested in the case where SU(2) is broken to U(1), where $g_{1}=g_{2}\neq g_{3}$ considered in the next section, however it’s easier to work out the beta functions for the fully anisotropic case first and then set $g_{1}=g_{2}$. Renormalization group beta-functions are generally prescription, or scheme, dependent. However for marginal perturbations, the one and two loop contributions are universal. [^5] To see this, consider a single coupling $g$ with $\beta_{g}=b_{2}g^{2}+b_{3}g^{3}+\ldots$ where $b_{2,3}$ are coefficients. The prescription dependence corresponds to a redefinition of the coupling $g^{\prime}=g^{\prime}(g)$. Let $g^{\prime}=g+cg^{2}+\ldots$. One easily sees that $\beta^{\prime}(g^{\prime})=b_{2}g^{\prime 2}+b_{3}g^{\prime 3}+\ldots.$ The implicit renormalization prescription of this article is based on the OPE (23), and doesn’t explicitly rely on any epsilon-expansion, nor Feynman diagrams for scalar fields. [^6] We do not address whether this OPE prescription can be shown to be equivalent to Feynman diagram perturbation theory and dimensional regularization to all orders in perturbation theory. However to 1-loop it was shown that these two-prescriptions do indeed agree to 1-loop for the closely related symplectic fermions LeClairNeubert . Referring to the title of Wilson-Fisher’s paper WilsonFisher , we work in precisely $4$ dimensions, not 3.993.99!
@@
where $\langle X\rangle_{0}$ is the unperturbed CFT result. The $n$-th term gives a contribution to the (n1)(n-1)-loop $\beta$ function.
III.1 1-loop
@@
III.2 2-loops
As we did above at 1-loop, at higher orders we use the OPE prescription to isolate $\log a$ divergences that contribute to the beta-functions. There are also $(\log a)^{n}$ divergences for $n>1$, which we interpret as arising from lower orders in perturbation theory which are already accounted for in the lower order beta-function. It turns out that the OPE (30) for the operators ${\cal O}^{a}$ is not sufficient to extract the appropriate $\log a$ contributions for 2-loops and beyond: one must split ${\cal O}^{a}$ into the product of $J^{a}$ and $\widetilde{J}^{a}$. It is useful to introduce diagrams that indicate the OPE structure that leads to a $\log a$ divergence of interest, with the building blocks shown in Figure 1.777Unlike Feynman diagrams, these diagrams are not a prescription for calculating the full correlation function, but are useful for isolating $\log a$ divergences. The 1-loop contribution is shown in Figure 2.
@@
The above result (41), based on Figure 3, is the only contribution at order $g^{3}$. For instance the diagram in Figure 4 gives a $\log^{2}a$ divergence which is already incorporated in the 1-loop beta function. The diagram in Figure 5 gives zero since it is proportional to $\delta^{ab}f^{abc}$. Figure 6 diagram could potentially lead to new terms $J^{a}(x)J^{a}(x)$ in the action rendering our original model non-renormalizable, however as is evident from the diagram, the two $J$’s are at different spacetime points, and furthermore leads to $\log^{2}a$ divergences.
@@
Figure 6: 2-loop diagram that potentially gives rise to additional JJJJ terms in the action.
III.3 3 loops
@@
$\beta_{g_{1}}=g_{2}g_{3}-\dfrac{\kappa}{4}g_{1}\left(g_{2}^{2}+g_{3}^{2}\right)+\dfrac{\kappa^{2}}{16}\left(g_{1}^{2}g_{2}g_{3}+g_{2}^{3}g_{3}+g_{3}^{3}g_{2}\right)$ (45)
plus cyclic permutations of $1,2,3$ for $\beta_{g_{2}}$ etc.
@@
IV SU(2) broken to U(1) and higher orders
IV.1 3 loops
@@ IV.2 Conjecture for all orders
To 3 loops, the above beta function (47) is identical to that for certain current-current perturbations in 2D reviewed in the Appendix. This implies that the algebraic structure of our OPE diagrams, i.e. it’s $f^{abc}$ and $\kappa$ dependence, is identical to the 2D current-current models. What is more surprising is that loop integrals $\int d^{4}x$ are quite different than for the 2D models, but nevertheless evaluate to give the same beta function. That this persists to 3-loops is already non-trivial, and there is no reason to expect this to fail at higher orders. For the 2D models, the nested geometric series were summed in Gerganov . Based on this we propose the following beta functions:
@@
One can easily check that to order $g^{4}$, the above agree exactly with the 3-loop results (47). What is not so obvious is that these beta functions preserve the RG invariant $Q$, which to all orders is now:
@@
The invariance condition (49) is exactly satisfied to all orders for this non-perturbative $Q$. The above beta functions do not constitute a gradient flow where $\beta_{g_{i}}=-\partial_{g_{i}}{\cal H}(g)$ for some height function ${\cal H}(g)$, since $\partial_{g_{1}}\beta_{g_{3}}\neq\partial_{g_{3}}\beta_{g_{1}}$. This is consistent with the existence of cyclic RG flows we find below since gradient flow is often associated with flows that are consistent with various forms of c-theorems.
@@
V Renormalization group trajectories
In this section we map out the RG flows in the various regimes of the couplings $g_{1},g_{3}$ for the case of SU(2) broken to U(1). We set $\kappa=1$ which is the correct value for our model. In the 1-loop approximation, the flows are shown in Figure 8. At 1-loop the RG invariant $Q=g_{1}^{2}-g_{3}^{2}$ implies the trajectories are hyperbolas. Along the separatrices $g_{3}=\pm g_{1}$ the SU(2) is unbroken. Here if $g_{3}>0$, the theory is asymptotically free in the UV where $g_{1}=g_{3}=0$. For $g_{3}<0$, the theory is marginally irrelevant and flows to strong coupling in the UV wherein there is no known UV fixed point. Just below the separatrices where SU(2) is broken to U(1) there is a line of fixed points $g_{1}=0$, which can be marginally relevant or irrelevant. Above the separatrices, there are no fixed points anywhere, and it was argued that this flow is actually cyclic ALDolls (see below). The cyclic flow is not at all rare for our model: small deviations from the SU(2) invariant flows along the diagonal separatrices can go either either way: either they flow to the line of fixed points or do not and are cyclic. The aim of this section is understand if this persists to higher orders based on the beta functions of the last section. We will base our analysis on the geometric re-summation of the OPE diagrams of the last section, which we argued leads to the all-orders beta functions (53). These beta functions are well defined in the limit $g\to\infty$, which is especially important for the cyclic flows which require $g_{3}$ flowing to $\pm\infty$.
@@ V Renormalization group trajectories
(i) $g_{3}>0$. Here the perturbation is relevant with $0<\Gamma<4$.
(ii) $4<g_{3}<0$. These are marginally irrelevant perturbations with $\Gamma>4$.
@@
[…]
@@
It was also shown in BLflow that it is consistent, and required, to endow the coupling constant space $(g_{1},g_{3})$ with the topology of a cylinder where one identifies $g_{3}=\pm\infty$, since $\beta(g_{1},g_{3})=\beta(g_{1},-g_{3})$ in the limit $|g_{3}|\to\infty$. This allows us to extend the flows to the entire $(g_{1},g_{3})$ plane, including through the poles at $g_{1}=\pm 4$, which are self-dual points under (57). We refer the reader to BLflow for more detailed explanations in the 2D case. This leads a variety of RG flows which we now itemize.
V.1 Massive Flows with fixed points
For case (i) above, namely $g_{3}>0$, the perturbation is relevant $0<\Gamma<4$ and flows to strong coupling in the IR. We interpret this as a massive phase where the IR fixed point is empty. In Figure 9 we plot the constant $Q$ RG trajectories based only on the 3 loop result (50), and one sees that the 1-loop flowchart in Figure 8 is not significantly modified. In comparison with the current-current perturbations reviewed in the Appendix, this is a sine-Gordon like phase.